OpenStudy (anonymous):

3x+4y=19 3x+6y=33 solve?

OpenStudy (anonymous):

Subtract... 3x + 6y = 33 3x + 4y = 19 -------------- 2y = 14 y = 7 --------------- 3x + 4y = 19 3x + 28 = 19 3x = -9 x = -3

OpenStudy (anonymous):

i feel lke i'm annoying Dx i just hate this stuff